# Fed-SB: A Silver Bullet for Extreme Communication Efficiency and Performance in (Private) Federated LoRA Fine-Tuning

### v1

@@ 3 Method
LoRA-SB for Fine-Tuning. LoRA-SB (Ponkshe et al., 2024) optimally approximates full FT gradients in low-rank spaces and demonstrates that its entire optimization trajectory aligns with the ideal low-rank projection of the full FT path. To achieve this, LoRA-SB fixes $\mathbf{A}$ and $\mathbf{B}$ while introducing a new trainable adapter $\mathbf{R}$ of size $r\times r$. Since $\mathbf{R}$ has rank $r$, it updates the pre-trained weight while maintaining rank $r$, making it highly parameter efficient. As a result, LoRA-SB consistently outperforms LoRA (and its variants) across various benchmarks while using 45–90x fewer trainable parameters.
@@
We show that Fed-SB effectively resolves all challenges in (private) federated FT while achieving state-of-the-art communication efficiency and performance. Table 1 highlights the advantages of Fed-SB over other methods.
Fed-SB: Exact Aggregation. Since only $\mathbf{R}$ is trainable, simple averaging of $\mathbf{R}$ across clients ensures exact aggregation without requiring updates to any other matrix. Further, the linearity of the global update with respect to the client-specific matrices $\mathbf{R}_{i}$ guarantees that exact aggregation occurs within rank $r$, preventing communication costs from scaling with the number of clients. This is because the server only needs to aggregate and transmit the $\mathbf{R}$ matrix. This can be proven directly by computing the global update $\Delta\mathbf{W}^{\text{agg}}$:
@@
$\Delta\mathbf{W}_{DP}-\Delta\mathbf{W}=\mathbf{B}\left(\mathbf{R}+\boldsymbol{\xi}_{B}\right)\mathbf{A}-\mathbf{B}\mathbf{R}\mathbf{A}\\ \implies\Delta\mathbf{W}_{DP}-\Delta\mathbf{W}=\mathbf{B}\boldsymbol{\xi}_{B}\mathbf{A}.$ (11)
@@
Fewer Learnable Parameters. The noise added to gradients for DP enforcement increases with the number of trainable parameters (Bassily et al., 2014; Abadi et al., 2016; Bun et al., 2014), potentially distorting learning and degrading performance. Reducing trainable parameters improves DP performance, provided the model retains sufficient task-specific expressivity.
@@
Lemma 1 establishes that reducing the number of learnable parameters enhances privacy guarantees under the same training setup. Specifically, achieving an equivalent level of privacy requires injecting less noise per parameter when fewer parameters are trained. Since LoRA-SB optimally approximates full fine-tuning gradients, its updates remain as effective as those in LoRA while benefiting from lower noise per update, resulting in a superior privacy-utility tradeoff. More generally, any reparameterization that reduces trainable parameters leads to a smaller accumulated privacy parameter $\epsilon$, thereby improving privacy performance—provided the reduction does not compromise learning.
Fed-SB: Pushing the Pareto Frontier. Fed-SB has significantly less communication costs than other federated FT methods. This is due to two key reasons: 1) LoRA-SB achieves performance comparable to or better than LoRA while requiring 45-90x fewer trainable parameters. 2) Fed-SB aggregates only the $r\times r$ trainable matrix $\mathbf{R}$, ensuring exact aggregation without additional communication overhead. This efficiency allows Fed-SB to leverage higher-rank updates without increasing communication costs. LoRA-SB typically operates at ranks 2–4x higher than LoRA, enabling Fed-SB to capture richer model updates. Retaining high-rank information is crucial in FL (Mahla et al., 2025) and is a key factor in the superior performance of FedEx-LoRA/FLoRA over FFA-LoRA and Fed-IT beyond just aggregation exactness. We empirically validate Fed-SB’s advantages through extensive experiments.
@@ 4 Experiments & Results
Method Rank # Comm. (M) ($\downarrow$) Accuracy ($\uparrow$)
BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Avg
@@
Table 2: Federated fine-tuning performance of Llama-3.2 3B across eight commonsense reasoning datasets. # Comm. denotes the number of parameters communicated per round. Best results are in bold.
Model Method Rank # Comm. (M) ($\downarrow$) Accuracy ($\uparrow$)
@@
Table 3: Federated fine-tuning performance of Mistral-7B and Gemma-2 9B on arithmetic reasoning benchmarks GSM8K and MATH. # Comm. denotes the number of parameters communicated per round. Best results are in bold.
Overview. We evaluate across three diverse NLP benchmarks, covering models that span from BERT-base (110M) to Gemma-2 (9B), thereby encompassing both masked and autoregressive architectures. Specifically, we fine-tune Mistral-7B (Jiang et al., 2023), Gemma-2 9B (Team et al., 2024), Llama-3.2 3B (Dubey et al., 2024), and BERT-base (Devlin, 2018). Experiments are conducted on a single NVIDIA A6000 GPU (48 GB), with results averaged over three random runs. Our experiments consider both performance and communication efficiency. To optimize memory efficiency, all base models (except BERT) are loaded in torch.bfloat16. Detailed dataset specifications can be found in Appendix D. For federated data distribution, we adopt a standard protocol where client datasets are randomly sampled, following established practice in FL (Sun et al., 2024; He et al., 2020; Lai et al., 2022). Detailed settings are provided in Appendix C.
Baselines. We evaluate Fed-SB against several SOTA federated FT approaches described previously, considering both private and non-private settings. Specifically, we compare it with FedIT, FedEx-LoRA, FLoRA, and FFA-LoRA. Where applicable, we also include comparisons with standard centralized LoRA (Hu et al., 2021).
@@ 4.1 Instruction Tuning
Details. We conduct experiments in the federated non-private setting across two reasoning tasks: commonsense reasoning and arithmetic reasoning. For commonsense reasoning, we fine-tune Llama-3.2 3B on CommonSense170K—a dataset aggregating eight commonsense reasoning corpora (Hu et al., 2023)—and evaluate its effectiveness across all constituent datasets. The experiments are performed in a cross-silo federated learning setup involving $5$ clients. For arithmetic reasoning, we fine-tune Mistral-7B (Jiang et al., 2023) and Gemma-2 9B (Team et al., 2024) on 20K samples from the MetaMathQA dataset (Yu et al., 2024) and assess their performance on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) benchmarks. In this setup, we distribute the federated training across $25$ clients. In both cases, we apply LoRA modules to the key, query, value, attention output, and all fully connected weights.
Results (Tables 2, 3). Our method significantly advances the performance-communication cost Pareto frontier, outperforming all previous baselines in both accuracy and communication efficiency across all models and benchmarks. Figure 3 further illustrates this significant improvement.
Method Rank # Params. (K) ($\downarrow$) Accuracy ($\uparrow$)
@@
Table 4: Centralized (Cent.) private fine-tuning of BERT-base on SNLI for varying values of $\epsilon$. A smaller $\epsilon$ indicates a stricter privacy budget. # Params. denotes the number of trainable parameters. Best results are in bold.
Method Rank # Comm. (K) ($\downarrow$) Accuracy ($\uparrow$)
@@
Table 5: Federated private fine-tuning of BERT-base on SNLI for varying values of $\epsilon$. A smaller $\epsilon$ indicates a stricter privacy budget. # Comm. denotes the number of parameters communicated per round. Best results are in bold.
Commonsense Reasoning (Table 2). Fed-SB ($r=200$) achieves an average accuracy improvement of 4.56% over FedIT while requiring lower communication cost. Additionally, Fed-SB ($r=200$) surpasses the previous SOTA performance methods FedEx-LoRA/FLoRA by 2.09%, while reducing communication cost by an impressive 31×. Notably, while the communication cost of FedEx-LoRA/FLoRA scales linearly with the number of clients, our method maintains a constant, client-independent communication cost. These results are obtained with just $5$ clients, implying that the full extent of our method’s communication efficiency is not fully depicted in this experiment. As the number of clients increases, the relative advantage of Fed-SB over existing methods grows even further.
Arithmetic Reasoning (Table 3): For Mistral-7B, Fed-SB ($r=200$) outperforms FedEx-LoRA/FLoRA on GSM8K by 1.90%, while achieving an impressive 234× reduction in communication cost. Additionally, Fed-SB ($r=200$) surpasses FFA-LoRA on GSM8K by 2.51%, with an approximately lower communication cost. Similarly, for Gemma-2 9B, Fed-SB ($r=200$) outperforms FedEx-LoRA/FLoRA on MATH by 0.86%, while reducing communication cost by 230×.
@@
Details. We fine-tune BERT-base (Devlin, 2018) on SNLI (Bowman et al., 2015), a standard benchmark for natural language inference. Following Hu et al. (2021), we apply LoRA modules only to the self-attention layers. Our evaluation considers two DP settings: a centralized private setup and a federated private setup. To enforce DP guarantees during training, we use the Opacus library (Yousefpour et al., 2021) with the DP-SGD optimizer (Abadi et al., 2016). In the federated setting, training is conducted in a cross-silo setup with $3$ clients. We conduct experiments across a range of privacy budgets, varying $\epsilon$ from $1$ to $10$.
Results (Tables 4, 5). Fed-SB consistently outperforms all prior baselines in both accuracy and communication/parameter efficiency across all privacy budgets in both settings. Figures 4, 5, and 6 further illustrate this significant improvement.
Centralized Private (Table 4). We first evaluate our method in the centralized setting against existing approaches. Fed-SB consistently outperforms previous methods while utilizing only a fraction of the parameters across all $\epsilon$ values. For instance, at $\epsilon=3$, Fed-SB ($r=64$) surpasses centralized LoRA and centralized FFA-LoRA by 9.28% and 2.05%, respectively, while using approximately 12x and 6x fewer parameters.
@@ 5 Conclusion
Existing LoRA-based federated FT methods either suffer from suboptimal updates or incur prohibitively high communication costs. We introduce Fed-SB, a federated adaptation of LoRA-SB that ensures exact aggregation while maintaining high communication efficiency. By training only a small $r\times r$ matrix and leveraging direct averaging, Fed-SB eliminates high-rank update costs and achieves communication efficiency independent of the number of clients. Fed-SB is particularly well-suited for private FT, as its linearity prevents noise amplification, and its reduced parameter count minimizes noise required for enforcing DP guarantees. It consistently outperforms existing methods across all models and tasks while reducing communication costs by up to 230x. These advantages establish Fed-SB as an efficient and scalable solution, setting a new Pareto frontier in (private) federated FT.
For future work, we plan to analyze the geometry of the $R$ matrices, as it is important for interpretability and may provide insights into the learning process. We also aim to study how the $R$ matrices change during training, which is key to understanding the fine-tuning process and characterizing its properties.
6 Limitations
While our approach is easily adaptable to other architectures, such as vision language models (VLMs) and vision transformers (VITs), we have not evaluated on such architectures. We have not yet extended our work to rank-heterogeneous settings, where clients operate with different ranks and computational budgets. Furthermore, we do not evaluate Fed-SB in scenarios with extreme data heterogeneity or significant class imbalances.
@@ Appendix A Proof of Lemma 1
The following result from Abadi et al. (2016) describes the relationship between noise variance, privacy parameters, number of optimization steps, batch size, and sample size in DP-SGD.
@@ Appendix B Related Work
Parameter-Efficient Fine-Tuning (PEFT). LoRA (Hu et al., 2021) has become ubiquitous for fine-tuning LLMs (Zhang et al., 2024a) by modeling weight updates as product of low-rank matrices. Several variants have been proposed to improve efficiency, stability, and adaptability. QLoRA (Dettmers et al., 2024) enables efficient fine-tuning through quantization strategies, reducing memory usage while maintaining performance. AdaLoRA (Zhang et al., 2023) dynamically allocates a layer-specific rank budget by assigning importance scores to individual weight matrices. LoRA-XS (Bałazy et al., 2024) further reduces trainable parameters by inserting a trainable matrix between frozen LoRA matrices. VeRA (Kopiczko et al., 2024) enhances parameter efficiency by learning shared adapters across layers. DoRA (Liu et al., 2024) decomposes the pre-trained matrix into two parts—magnitude and direction—and applies LoRA modules only to the direction component. PiSSA (Meng et al., 2024) improves adaptation by initializing adapters using the singular value decomposition (SVD) of pre-trained weights. rsLoRA (Kalajdzievski, 2023) introduces a rank-scaling factor to stabilize learning. LoRA-SB (Ponkshe et al., 2024) provably approximates gradients optimally in low-rank spaces, achieving superior performance with significantly higher parameter efficiency.
@@
Federated Fine-Tuning. Federated Learning (FL) consists of a centralized global model and multiple clients, each with its own local dataset and computational capacity. The global model is updated by aggregating client updates (Kairouz et al., 2021). FedBERT (Tian et al., 2022) focuses on federated pre-training, while other methods work on federated fine-tuning (Zhang et al., 2022; Kuang et al., 2024; Babakniya et al., 2023). Fed-IT (Zhang et al., 2024c) aggregates low-rank adapters across clients using standard federated averaging (McMahan et al., 2017) before updating the global model. To address inexact aggregation, FedEx-LoRA (Singhal et al., 2025) introduces an error matrix to correct residual errors, ensuring more precise updates. FLoRA (Wang et al., 2024) follows the same exact aggregation principle by stacking matrices and extends this approach to heterogeneous rank settings. FFA-LoRA (Sun et al., 2024) mitigates aggregation inexactness by freezing $\mathbf{A}$ and updating only the trainable low-rank adapter, averaging the latter to compute the global update. In some scenarios, clients require heterogeneous LoRA ranks due to varying computational budgets (Zhao et al., 2018; Li et al., 2019). Methods like HetLoRA (Cho et al., 2024) enable rank heterogeneity through self-pruning and sparsity-aware aggregation strategies, but incur significant overhead.

### v2

@@ 3 Method
LoRA-SB for Fine-Tuning. LoRA-SB (39) optimally approximates full FT gradients in low-rank spaces and demonstrates that its entire optimization trajectory aligns with the ideal low-rank projection of the full FT path. To achieve this, LoRA-SB fixes $\mathbf{A}$ and $\mathbf{B}$ while introducing a new trainable adapter $\mathbf{R}$ of size $r\times r$. Since $\mathbf{R}$ has rank $r$, it updates the pre-trained weight while maintaining rank $r$, making it highly parameter efficient. As a result, LoRA-SB consistently outperforms LoRA (and variants) across benchmarks while using 45–90x fewer trainable parameters.
@@
We show that Fed-SB effectively resolves all challenges in (private) federated FT while achieving state-of-the-art communication efficiency and performance. Table 1 highlights the advantages of Fed-SB over other methods. Since Fed-SB fixes the adapter matrices $A$ and $B$ throughout training, their initialization is crucial for effective learning. We adopt the update-based initialization strategy from LoRA-SB, which we detail in Appendix C.
Fed-SB: Exact Aggregation. Since only $\mathbf{R}$ is trainable, simple averaging of $\mathbf{R}$ across clients ensures exact aggregation without any updates to any other matrix. Further, the linearity of the global update with respect to the client-specific matrices $\mathbf{R}_{i}$ guarantees that exact aggregation occurs within rank $r$, preventing communication costs from scaling with number of clients. This is because the server only needs to aggregate and transmit $\mathbf{R}$, which can be proven by computing the global update $\Delta\mathbf{W}^{\text{agg}}$:
@@
$\Delta\mathbf{W}_{DP}-\Delta\mathbf{W}=\mathbf{B}\left(\mathbf{R}+\boldsymbol{\xi}_{B}\right)\mathbf{A}-\mathbf{B}\mathbf{R}\mathbf{A}\implies\Delta\mathbf{W}_{DP}-\Delta\mathbf{W}=\mathbf{B}\boldsymbol{\xi}_{B}\mathbf{A}.$ (13)
@@
Fewer Learnable Parameters. The noise added to gradients for DP enforcement increases with the number of trainable parameters (4, 1, 9), potentially distorting learning and degrading performance. Reducing trainable parameters improves DP performance, provided the model retains sufficient task-specific expressivity.
@@
Lemma 1 establishes that reducing the number of learnable parameters enhances privacy guarantees under the same training setup. Specifically, achieving an equivalent level of privacy requires injecting less noise per parameter when fewer parameters are trained. Since LoRA-SB optimally approximates full fine-tuning gradients, its updates remain as effective as those in LoRA while benefiting from lower noise per update, resulting in a superior privacy-utility tradeoff. More generally, any reparameterization that reduces trainable parameters leads to a smaller accumulated privacy parameter $\epsilon$, thereby improving performance, provided the reduction does not compromise learning.
Fed-SB: Pushing the Pareto Frontier. Fed-SB has significantly less communication costs than other federated FT methods. This is due to two key reasons: 1) LoRA-SB achieves performance comparable to or better than LoRA while requiring 45-90x fewer trainable parameters. 2) Fed-SB aggregates only the $r\times r$ trainable matrix $\mathbf{R}$, ensuring exact aggregation without additional communication overhead. This allows Fed-SB to leverage higher-rank updates without increasing communication costs. LoRA-SB typically operates at ranks 2-4x higher than LoRA, enabling Fed-SB to capture richer updates. Retaining high-rank information is crucial in FL (34) and a key factor in the superior performance of FedEx-LoRA/FLoRA over FFA-LoRA/Fed-IT beyond just aggregation exactness.
While our main focus is on the rank-homogeneous setting (where all clients use the same adapter rank), we also extend Fed-SB to support rank-heterogeneous clients, where each client trains with its own local rank budget. Additional details and results are provided in Table 7 (Appendix D), where we show that the rank-heterogeneous setup achieves performance comparable to the homogeneous rank settings.
@@ 4 Experiments & Results
Table 2: Federated fine-tuning of Llama-3.2 3B across eight commonsense reasoning datasets. # Comm. denotes the number of parameters communicated per round (in M). Best results are in bold.
Method Rank # Comm. ($\downarrow$) Accuracy ($\uparrow$)
BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Avg.
@@
Table 3: Federated fine‑tuning of Llama‑3.2 3B across eight commonsense reasoning datasets, in a highly data‑heterogeneous setting, where each client is trained on a distinct dataset. # Comm. denotes the number of parameters communicated per round (in M). Best results are in bold.
Method Rank # Comm. ($\downarrow$) Accuracy ($\uparrow$)
BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Avg.
FedIT 32 48.63 60.89 78.22 69.92 73.18 67.88 81.21 67.04 66.91 70.80
FFA-LoRA 32 24.31 60.73 76.91 65.37 65.18 61.89 79.41 62.92 63.12 67.17
FedEx-LoRA 32 243.15 62.55 79.36 71.41 78.12 72.45 82.89 67.88 70.25 73.13
FLoRA 32 243.15 62.55 79.36 71.41 78.12 72.45 82.89 67.88 70.25 73.13
Fed-SB 120 2.83 61.41 78.13 71.02 78.24 71.78 82.45 67.12 68.83 72.65
Fed-SB 160 5.02 62.34 79.05 72.39 80.52 74.67 83.18 68.64 70.12 73.98
Fed-SB 200 7.85 63.28 80.34 73.56 82.07 76.01 84.01 69.02 72.46 75.21
Table 4: Federated fine-tuning of Mistral-7B and Gemma-2 9B on GSM8K and MATH. # Comm. denotes the number of parameters communicated per round (in M). Best results are in bold.
Model Method Rank # Comm. ($\downarrow$) Accuracy ($\uparrow$)
@@
Overview. We evaluate across three diverse NLP benchmarks, covering models that span from BERT-base (110M) to Gemma-2 (9B), thereby encompassing both masked and autoregressive architectures. Specifically, we fine-tune Mistral-7B (23), Gemma-2 9B (47), Llama-3.2 3B (16), and BERT-base (15). Our experiments consider both performance and communication efficiency. Detailed experimental and dataset specifications are provided in Appendix G and H, respectively. For federated data distribution, we adopt a standard protocol where client datasets are randomly sampled, following established practice in FL (44, 19, 29). We conduct experiments on a single NVIDIA A6000 GPU (48 GB) and report the average results from three independent runs.
Baselines. We evaluate against several SOTA federated FT approaches described previously, considering both private and non-private settings. Specifically, we compare it with FedIT, FedEx-LoRA, FLoRA, and FFA-LoRA. Where applicable, we also include comparisons with standard LoRA (21).
@@ 4.1 Instruction Tuning
Details. We conduct experiments in the federated non-private setting across two reasoning tasks: commonsense reasoning and arithmetic reasoning. For commonsense reasoning, we fine-tune Llama-3.2 3B on CommonSense170K, a dataset aggregating eight commonsense reasoning corpora (22), and evaluate its effectiveness across all constituent datasets. The experiments are performed in a cross-silo federated learning setup involving $5$ clients.
We also evaluate Fed-SB under extreme data heterogeneity. Instead of randomly sampling examples for each client, we assign each constituent dataset to a distinct client, resulting in a highly non-IID 8-client setup. Each client trains on a distinct distribution, with varying dataset sizes.
For arithmetic reasoning, we fine-tune Mistral-7B (23) and Gemma-2 9B (47) on 20K samples from the MetaMathQA dataset (55) and assess their performance on the GSM8K (13) and MATH (20) benchmarks. In this setup, we distribute the federated training across $25$ clients. In both cases, we apply LoRA modules to the key, query, value, attention output, and all fully connected weights.
Results (Tables 2, 3, 4). Our method achieves state-of-the-art performance, outperforming all previous baselines in both accuracy and communication efficiency across all models and benchmarks. Figure 3 further illustrates this significant improvement. Additional results on the effect of varying rank are reported in Table 8 in Appendix E.
Table 5: Centralized (Cent.) private fine-tuning of BERT-base on SNLI for varying values of $\epsilon$. A smaller $\epsilon$ indicates a stricter privacy budget. # Params. denotes the number of trainable parameters (in K). Best results are in bold.
Method Rank # Params. ($\downarrow$) Accuracy ($\uparrow$)
@@
Table 6: Federated private fine-tuning of BERT-base on SNLI for varying values of $\epsilon$. A smaller $\epsilon$ indicates a stricter privacy budget. # Comm. denotes the number of parameters communicated per round (in K). Best results are in bold.
Method Rank # Comm. ($\downarrow$) Accuracy ($\uparrow$)
@@
Commonsense Reasoning (Table 2). Fed-SB ($r=200$) achieves an average improvement of 4.56% over FedIT while requiring lower communication cost. Additionally, Fed-SB ($r=200$) surpasses the previous SOTA performance methods FedEx-LoRA/FLoRA by 2.09%, while reducing communication cost by an impressive 31×. Notably, while the communication cost of FedEx-LoRA/FLoRA scales linearly with the number of clients, our method maintains a constant, client-independent communication cost. These results are obtained with just $5$ clients, implying that the full extent of our method’s communication efficiency is not fully depicted here. As the number of clients increases, the relative advantage of Fed-SB over existing methods grows even further.
Highly Data-Heterogenous Setting (Table 3). Fed-SB significantly outperforms all other methods even in this highly non-IID setting. Specifically, Fed-SB ($r=200$) surpasses the previous state-of-the-art methods, FedEx-LoRA and FLoRA, by 2.08% in accuracy while achieving a remarkable 31× reduction in communication cost.
Arithmetic Reasoning (Table 4). For Mistral-7B, Fed-SB ($r=200$) outperforms FedEx-LoRA/FLoRA on GSM8K by 1.90%, while achieving an impressive 234× reduction in communication cost. Additionally, Fed-SB ($r=200$) surpasses FFA-LoRA on GSM8K by 2.51%, with approximately lower communication cost. For Gemma-2 9B, Fed-SB ($r=200$) outperforms FedEx-LoRA/FLoRA on MATH by 0.86%, while reducing communication cost by 230×.
@@
Details. We fine-tune BERT-base (15) on SNLI (8), a standard benchmark for natural language inference. Following LoRA(21), we apply LoRA modules only to the self-attention layers. Our evaluation considers two DP settings: a centralized private setup and a federated private setup. To enforce DP guarantees during training, we use the Opacus library (53) with the DP-SGD optimizer (1). In the federated setting, training is conducted in a cross-silo setup with $3$ clients. We conduct experiments across a range of privacy budgets, varying $\epsilon$ from $1$ to $10$.
Results (Tables 5, 6). Fed-SB consistently outperforms all prior baselines in both accuracy and communication/parameter efficiency across all privacy budgets in both settings. Figures 4, 5, and 6 further illustrate this significant improvement. Further experiments analyzing the impact of rank variation are given in Table 9 (Appendix E).
Centralized Private (Table 5). Fed-SB showcases significant improvement over other methods while using only a fraction of the parameters, across all $\epsilon$ values. For instance, at $\epsilon=3$, Fed-SB ($r=64$) surpasses centralized LoRA and centralized FFA-LoRA by 9.28% and 2.05%, respectively, while using $\approx$ 12x and 6x fewer parameters.
@@
4.3 Memory and Training Time
Memory. Fed-SB needs lower per-client training memory relative to all other baselines by substantially reducing the number of trainable parameters. Notably, this advantage holds even when Fed-SB is trained with a higher rank ($r=200$), where it still requires less memory than competing methods at a lower rank ($r=32$). We note that the peak memory usage of Fed-SB never exceeds that of any other federated LoRA-based baseline. Detailed analysis is provided in Table 10 (Appendix F).
Training Time. Fed-SB introduces a negligible training time overhead ($\approx 2\%$) relative to other methods, attributable to its initialization step. We benchmark this overhead in Table 11 (Appendix F).
@@ 5 Conclusion
Existing LoRA-based federated FT methods either suffer from suboptimal updates or incur prohibitively high communication costs. We introduce Fed-SB, a federated adaptation of LoRA-SB that ensures exact aggregation while maintaining high communication efficiency. By training only a small $r\times r$ matrix and leveraging direct averaging, Fed-SB eliminates high-rank update costs and achieves communication efficiency independent of the number of clients. Fed-SB is particularly well-suited for private FT, as its linearity prevents noise amplification, and its reduced parameter count minimizes noise required for enforcing DP guarantees. It consistently achieves a new state-of-the-art across all models and tasks while reducing communication costs by up to 230x. These advantages establish Fed-SB as an efficient and scalable solution for (private) federated FT.
@@ Appendix A Proof of Lemma 1
The following result (1) describes the relationship between noise variance, privacy parameters, number of optimization steps, batch size, and sample size in DP-SGD.
@@ Appendix B Related Work
Parameter-Efficient Fine-Tuning (PEFT). LoRA (21) has become ubiquitous for fine-tuning LLMs (57) by modeling weight updates as product of low-rank matrices. Several variants have been proposed to improve efficiency, stability, and adaptability. QLoRA (14) enables efficient fine-tuning through quantization strategies, reducing memory usage while maintaining performance. AdaLoRA (60) dynamically allocates a layer-specific rank budget by assigning importance scores to individual weight matrices. LoRA-XS (5) further reduces trainable parameters by inserting a trainable matrix between frozen LoRA matrices. VeRA (27) enhances parameter efficiency by learning shared adapters across layers. DoRA (32) decomposes the pre-trained matrix into two parts—magnitude and direction—and applies LoRA modules only to the direction component. PiSSA (36) improves adaptation by initializing adapters using the singular value decomposition (SVD) of pre-trained weights. rsLoRA (25) introduces a rank-scaling factor to stabilize learning. LoRA-SB (39) provably approximates gradients optimally in low-rank spaces, achieving superior performance with significantly higher parameter efficiency.
@@
Federated Fine-Tuning. Federated Learning (FL) consists of a centralized global model and multiple clients, each with its own local dataset and computational capacity. The global model is updated by aggregating client updates (24). FedBERT (48) focuses on federated pre-training, while other methods work on federated fine-tuning (61, 28, 3). Fed-IT (59) aggregates low-rank adapters across clients using standard federated averaging (35) before updating the global model. To address inexact aggregation, FedEx-LoRA (43) introduces an error matrix to correct residual errors, ensuring more precise updates. FLoRA (52) follows the same exact aggregation principle by stacking matrices and extends this approach to heterogeneous rank settings. FFA-LoRA (44) mitigates aggregation inexactness by freezing $\mathbf{A}$ and updating only the trainable low-rank adapter, averaging the latter to compute the global update. In some scenarios, clients require heterogeneous LoRA ranks due to varying computational budgets (62, 30). Methods like HetLoRA (10) enable rank heterogeneity through self-pruning and sparsity-aware aggregation strategies, but incur significant overhead.